# CO-VADA: A Confidence-Oriented Voice Augmentation Debiasing Approach for Fair Speech Emotion Recognition

### v1

@@ VII-E SER Performance and Fairness Across Emotion Categories
We further evaluated the SER performance and fairness of the CO-VADA across different emotion categories. As shown in Table VI, CO-VADA consistently reduced gender-based disparities while maintaining competitive classification performance across all categories. These results demonstrate the robustness of CO-VADA in handling diverse emotional expressions.
Table VI: SER Performance and gender fairness across emotion categories.
Emotion Method Macro F1 $\uparrow$ $\boldsymbol{\mathrm{TPR}_{\text{gap}}}$downarrow $\boldsymbol{\mathrm{DP}_{\text{gap}}}$downarrow
Angry ERM 0.542 0.277 0.097
CO-VADA 0.534 0.270 0.104
Happy ERM 0.535 0.379 0.134
CO-VADA 0.540 0.374 0.118
Sad ERM 0.541 0.240 0.114
CO-VADA 0.536 0.229 0.090
Fear ERM 0.594 0.279 0.130
CO-VADA 0.594 0.229 0.103
Neutral ERM 0.611 0.293 0.105
CO-VADA 0.606 0.247 0.080
Disgust ERM 0.577 0.273 0.099
CO-VADA 0.576 0.224 0.081
VIII Conclusion and Future Work
@@
While our method is not limited to gender bias, due to the limitations of the corpus and its labels, we evaluated the proposed approach only on a dataset that contains gender bias. Future work may explore other sources of bias, such as semantic [61], linguistic [62], or stylistic variation [63]. Alternative strategies that manipulate latent representations instead of generating audio could achieve more efficient and flexible debiasing.

### v2

@@ VII-E SER Performance and Fairness Across Emotion Categories
We further evaluated the SER performance and fairness of CO-VADA across different emotion categories. As shown in Table VI, we report the original results of CO-VADA for each emotion class, with subscripts indicating the relative changes compared to the corresponding ERM baseline for that class. That is, each comparison uses the ERM result specific to the same emotion category, rather than the overall ERM average reported earlier.
The proposed CO-VADA consistently reduced gender-based disparities across emotions, while maintaining competitive classification performance in most categories. These results demonstrate the robustness of CO-VADA in handling diverse emotional expressions.
Table VI: SER performance and gender fairness of CO-VADA across emotion categories. Subscripts indicate relative changes compared to the ERM baseline.
Emotion F1 $\uparrow$ TPRgap $\downarrow$ DPgap $\downarrow$
Angry 0.534 -1.48% 0.270 -2.53% 0.104 +7.22%
Happy 0.540 +0.93% 0.374 -1.32% 0.118 -11.94%
Sad 0.536 -0.92% 0.229 -4.58% 0.090 -21.05%
Fear 0.594 0.00% 0.229 -17.92% 0.103 -20.77%
Neutral 0.606 -0.82% 0.247 -15.71% 0.080 -23.81%
Disgust 0.576 -0.17% 0.224 -17.58% 0.081 -18.18%
VIII Conclusion, Limitations, and Future Work
@@
A key limitation of our approach is the reliance on prediction confidence to separate bias-guiding and bias-contrary samples. While effective in practice, this heuristic may conflate underrepresented but easy samples with overrepresented but difficult ones.
Moreover, while our method is not limited to gender bias, due to the limitations of the corpus and its labels, we evaluated the proposed approach only on a dataset that contains gender bias. Future work may explore other sources of bias, such as semantic [61], linguistic [62], or stylistic variation [63]. Besides, investigate alternative debiasing strategies that operate in the latent space or directly manipulate embeddings potentially enabling more efficient and flexible mitigation without relying on audio synthesis.